# Learning System Dynamics without Forgetting

### v1

@@ Abstract
Predicting the trajectories of systems with unknown dynamics (i.e. the governing rules) is crucial in various research fields, including physics and biology. This challenge has gathered significant attention from diverse communities. Most existing works focus on learning fixed system dynamics within one single system. However, real-world applications often involve multiple systems with different types of dynamics or evolving systems with non-stationary dynamics (dynamics shifts). When data from those systems are continuously collected and sequentially fed to machine learning models for training, these models tend to be biased toward the most recently learned dynamics, leading to catastrophic forgetting of previously observed/learned system dynamics. To this end, we aim to learn system dynamics via continual learning. Specifically, we present a novel framework of Mode-switching Graph ODE (MS-GODE), which can continually learn varying dynamics and encode the system-specific dynamics into binary masks over the model parameters. During the inference stage, the model can select the most confident mask based on the observational data to identify the system and predict future trajectories accordingly. Empirically, we systematically investigate the task configurations and compare the proposed MS-GODE with state-of-the-art techniques. More importantly, we construct a novel benchmark of biological dynamic systems, featuring diverse systems with disparate dynamics and significantly enriching the research field of machine learning for dynamic systems.
Figure 1: Illustration of one of the biological cellular systems involved in our work: RAN-regulated nucleocytoplasmic transport [1]. Briefly speaking, this model of translocation of a cargo protein via the nuclear pore is based on the Ran protein. In such systems, multiple factors could change and alter the entire dynamics.
@@ 1 Introduction
Scientific research often involves systems composed of interacting objects, such as multi-body systems in physics and cellular systems in biology, with their evolution governed by underlying dynamic rules. However, due to potentially unknown or incomplete dynamic rules or incomplete observations, deriving explicit equations to simulate system evolution can be extremely challenging. As a result, data-driven approaches based on machine learning have emerged as a promising solution for predicting the future trajectories of system states purely from observational data. For instance, the Interaction Network (IN) model [2] explicitly learns interactions between pairs of objects and has demonstrated superior performance in simulated physics systems, showcasing the potential of machine learning in studying physical system dynamics. IN has inspired many subsequent works [3, 4, 5, 6]. Later, to enable the modeling of incomplete and temporarily irregular system observations, ODE-based models [7, 8] were proposed to learn the continuous dynamics of the systems.
Despite the success of these methods, existing approaches typically assume the fixed type of dynamics within the target system. However, in many real-world scenarios, system dynamics are subject to change over time, or multiple systems may involve different types of dynamics. When data from those systems are continuously collected and sequentially fed to machine learning models for training, an ideal model is expected to continually learn varying dynamics without forgetting the previously observed ones and correctly infer any type of underlying dynamics during the test stage. For example, in cellular dynamics modeling (Figure 1), the dynamics of the same set of variables are subject to change when the kinetic factors change. Besides, a widely adopted task in dynamic system learning is predicting the trajectories of objects in an n-body system based on the observed trajectories [7, 2]. As illustrated in Figure 2, systems may be governed by different dynamics including various interaction types (e.g., elastic force or electrostatic force) and different interaction strengths (e.g., different amounts of charges on charged objects). In such scenarios, models that continually learn from systems with varying dynamics may encounter the catastrophic forgetting problem. In other words, the model can only accurately predict the most recently observed system dynamics while failing to predict earlier system dynamics due to forgetting.
@@
Targeting the challenge, we propose a novel mode-switching graph ODE (MS-GODE) model, which can continually learn varying system dynamics and automatically switch to the corresponding mode during the test stage. Overall, MS-GODE follows the Variational AutoEncoder (VAE) [9] structure. Given the observational data, an encoder network first encodes the data into latent states, upon which an ODE-based generator predicts the future system trajectories within the latent space. Finally, a decoder network maps the predicted latent states back into the data space. This framework can effectively learn the continuous system dynamics and is suitable for system observations with irregular temporal sampling intervals, such as the biological systems studied in our work. However, for continual learning over different types of system dynamics, this framework alone is prone to catastrophic forgetting, as the new dynamic patterns may overwrite old knowledge encoded in the model parameters. Inspired by the advancements of sub-networks [10, 11, 12], we fix the model weights and optimize a unique binary mask over the parameters for each system during training. Unlike standard model training, which encodes learned patterns solely in model weights, each type of dynamics in MS-GODE is encoded in the topology of a sub-network (the binary mask), and the backbone with fixed weight values. In the test stage, the most suitable mask, which can most accurately reconstruct the given observation, is chosen and applied to the backbone network to form a system specific sub-network. This approach avoids catastrophic forgetting because the backbone network parameters are fixed, and system-specific knowledge is encoded in separate binary masks.
@@
Besides the technical contribution, we have also created a novel dynamic system benchmark consisting of biological cellular systems based on the VCell platform [13, 14, 15]. Compared to the widely adopted simulated physics systems, cellular systems contain heterogeneous objects and interactions, offering richer and more challenging patterns of system dynamics. Therefore, the novel benchmark will significantly enhance the research of machine learning-based system dynamics prediction.
In experiments, we thoroughly investigate how the system sequence configuration influences the model performance when learning sequentially and then empirically demonstrate that our proposed MS-GODE outperforms existing state-of-the-art techniques.
@@ 2.1 Learning based dynamic System Prediction
In recent years, neural networks, especially graph neural networks (GNNs), have been proved to be promising in predicting the complex evolution of systems consisting of interacting objects [2, 3, 4, 5, 6]. This was firstly demonstrated by [2] with Interaction Network (IN), which iteratively infers the effects of the pair-wise interactions within a system and predicts the changes of the system states. Following IN, [3] proposed neural relational inference (NRI) to predict systems consisting of objects with unknown relationship. [16] proposed Hierarchical Relation Network (HRN) that extends the predictions to systems consisting of deformable objects. [4] proposed Hamiltonian ODE graph network (HOGN), which injects Hamiltonian mechanics into the model as a physically informed inductive bias. Later, to better consider the intrinsic symmetry of the target systems, GNNs with different invariance and equivariane are proposed [17, 5, 18, 19, 20, 6].
Considering that the observation of real-world systems may be incomplete and irregularly samples, [7] proposed LG-ODE, which is capable of generating continuous system dynamics based on the latent ordinary differential equations. Following LG-ODE, [8] proposed Coupled Graph ODE (CG-ODE) to further extend the continuous dynamics prediction to systems with unknown relationship. Similar ideas are also adopted in other time series research [21]. Our work is closely related to the ODE based models [21, 7, 8], and we also encode the system dynamics within the irregularly and partially observed data via spatial-temporal GNN structures [7, 8]. Although these above mentioned methods have made significant contributions to dynamic system prediction from different perspectives, they have been restricted to learning a single system with fixed dynamics.
@@ 2.2 Continual Learning & Masked Networks
Existing continual learning methods can be briefly categorized into three types [22, 23, 24, 25]. Regularization based methods slow down the adaption of important model parameters via regularization terms, so that the forgetting problem is alleviated [26, 27]. For example, Elastic Weight Consolidation (EWC) [28] and Memory Aware Synapses (MAS) [29] estimate the importance of the model parameters to the learned tasks, and add penalty terms to slow down the update rate of the parameters that are important to the previously learned tasks. Second, experience replay based methods replay the representative data stored from previous tasks to the model when learning new tasks to prevent forgetting [30, 31, 32, 33]. For example, Gradient Episodic Memory (GEM) [34] leverages the gradients computed based on the buffered data to modify the gradients for learning the current task and avoid the negative interference between learning different tasks. Finally, parameter isolation based methods gradually introduce new parameters to the model for new tasks to prevent the parameters that are important to previous tasks [35, 36]. For example, Progressive Neural Network (PNN) [37] allocate new network branches for new tasks, such that the learning on new tasks does not modify the parameters encoding knowledge of the old tasks. Our proposed MS-GODE also belongs to the parameter-isolation based methods, and is related to Supermasks in Superposition (SupSup) [12] and the edge-popup algorithm [11]. SupSup studies continual learning for classification tasks with an output entropy based mask selection, which is not applicable to our task. Edge-popup algorithm provides a simple yet efficient strategy to select a sub-network from the original network, and is adopted by us to optimize the binary masks over the model parameters. As far as we are concerned, existing continual learning works have not been applied to the dynamic system prediction targeted by this paper. Therefore, by customizing the representative continual learning techniques to our task as baselines, we also contribute to the community by extending the applicable areas of the existing methods and by demonstrating the advantages and disadvantages of different methods on a new task.
@@ 3.1 Preliminaries
In our work, the model is required to sequentially learn on multiple systems with different dynamics. Each system is composed of multiple interacting objects, therefore is naturally represented as a graph $\mathcal{G}=\{\mathbb{V},\mathbb{E}\}$. $\{\mathbb{V}$ is the node set denoting the objects of the systems, and $\mathbb{E}$ is the edge set containing the information of the relation and interaction between the objects. Based on the edge set $\mathbb{E}$, the spatial neighborhood of a node is defined as $\mathcal{N}_{s}(v)=\{u|e_{u,v}\in\mathbb{E}\}$. Each object node $v$ is accompanied with observational data containing the observed states at certain time steps $\mathbb{X}_{v}=\{\mathbf{x}_{v}^{t}|t\in\mathbb{T}_{v}\}$. With the system structured as a graph, the observation of the system evolution trajectory is naturally a spatial temporal graph, of which each node is an observed state $\mathbf{x}_{v}^{t}$. In the multi-body system example, the observation is the trajectories of the particles over time. In the following, we will refer to $\mathbf{x}_{v}^{t}$) as a ‘state’ for convenience. The set $\mathbb{T}_{v}$ contains the time steps (real numbers) when the states of $v$ are observed and can vary across different objects. For the prediction task, the observations lie within a certain period, i.e. $\bigcup_{v\in\mathbb{V}}\mathbb{T}_{v}\in$[$t_{0},t_{1}$], and the task is to predict the system states at future time steps beyond $t_{1}$. We denote the future time steps to predict for an object $v$ as $\mathbb{T}_{v}^{pred}$. In our continual dynamics learning setting, the training stage requires a model to sequentially learn on multiple systems with different dynamics and potentially different objects (i.e. $\mathbb{V}$).
Figure 3: The upper part illustrates the workflow of MS-GODE during mask selection and inference, which are denoted by dashed and solid lines. During mask selection, the observation is split into two parts, and the first part is fed into the model for selecting the mask that can best predict the second part (simulated prediction). During inference, the entire observation is fed into the model to predict the unknown future states. The lower part demonstrates the structure of the masked encoder, masked generator, and masked decoder. Different components fetch their corresponding mask from the mask pool and apply the mask onto the parameters.
@@ 3.2 Framework Overview
In this subsection, we provide a high-level introduction on the workflow of MS-GODE (Figure 3), while the details of each component are provided in the following subsections.
Overall, MS-GODE consists of three core components: 1) An encoder network $\mathrm{Enc}(\cdot;\theta_{E})$ parameterized by the parameters $\theta_{E}$ for encoding the observational data into the latent space; 2) A ODE-based generator $\mathrm{Gen}(\cdot;\theta_{G})$ parameterized by $\theta_{G}$ for predicting the future states of the system within the latent space; 3) A decoder network $\mathrm{Dec}(\cdot;\theta_{D})$ parameterized by $\theta_{D}$ for transforming the predicted latent states back into the data space. Within a standard learning scheme, the model will be trained and optimized by updating the parameters $\{\theta_{E},\theta_{G},\theta_{D}\}$. However, as introduced above, when continually training the model on multiple systems with different dynamics, this learning scheme will bias the model towards the most recently observed system, causing the catastrophic forgetting problem. In this work, inspired by the recent advances in sub-network learning [11], we propose to avoid directly optimizing the parameters $\{\theta_{E},\theta_{G},\theta_{D}\}$. Instead, we train the model by optimizing the connection topology of the backbone model and encode the dynamics of each system into a sub-network. This is equivalent to encoding each system-specific dynamics pattern into a binary mask overlaying the shared parameters with fixed values. In this way, the interference between learning on different systems with different dynamics, i.e. the cause of the forgetting problem, can be avoided. Moreover, this approach is also memory efficient since the space for storing the binary masks is negligible. With the system-specific masks, the three model components are formulated as:

### v2

@@ Abstract
Observation-based trajectory prediction for systems with unknown dynamics is essential in fields such as physics and biology. Most existing approaches are limited to learning within a single system with fixed dynamics patterns. However, many real-world applications require learning across systems with evolving dynamics patterns, a challenge that has been largely overlooked. To address this, we systematically investigate the problem of Continual Dynamics Learning (CDL), examining task configurations and evaluating the applicability of existing techniques, while identifying key challenges. In response, we propose the Mode-switching Graph ODE (MS-GODE) model, which integrates the strengths LG-ODE and sub-network learning with a mode-switching module, enabling efficient learning over varying dynamics. Moreover, we construct a novel benchmark of biological dynamic systems for CDL, Bio-CDL, featuring diverse systems with disparate dynamics and significantly enriching the research field of machine learning for dynamic systems. Our code available at https://github.com/QueuQ/MS-GODE.
Figure 1: Illustration of the key components of one biological cellular system studied in our work: the RAN-regulated nucleocytoplasmic transport (Moore, 2013). Briefly speaking, this model depicts the translocation of cargo proteins (Exportin) via nuclear pores with the assistance of RAN proteins. RAN is first activated (RAN*) and then binds to cargo forming a complex. Next, the complex is translocated across the nuclear membrane into the cytoplasm with the assistance of RAN. Finally, RAN and Exportin are dissociated.
@@ 1 Introduction
Scientific research often involves systems composed of interacting objects, such as multi-body systems in physics and cellular systems in biology, with their evolution governed by underlying dynamic rules. However, due to potentially unknown or incomplete dynamic rules or incomplete observations, deriving explicit equations to simulate system evolution can be extremely challenging. As a result, data-driven approaches based on machine learning have emerged as a promising solution for predicting the future trajectories of system states purely from observational data. For instance, the Interaction Network (IN) model (Battaglia et al., 2016) explicitly learns interactions between pairs of objects and has demonstrated superior performance in simulated physics systems, showcasing the potential of machine learning in studying physical system dynamics. IN has inspired many subsequent works (Kipf et al., 2018; Sanchez-Gonzalez et al., 2019; Huang et al., 2022; Liu et al., 2024). Later, to enable the modeling of incomplete and temporarily irregular system observations, ODE-based models (Huang et al., 2020; 2021) were proposed to learn the continuous dynamics of the systems.
Despite the success of these methods, existing approaches are often limited to learning within a single system with a fixed type of dynamics. However, in many real-world scenarios, system dynamics are subject to change over time. For example, in cellular systems (Figure 1), the dynamics of a set of variables are subject to change when the kinetic factors are altered. Besides, as a widely adopted task in the field, predicting the trajectories of n-body physical systems (Huang et al., 2020; Battaglia et al., 2016) may also require a model to learn over systems governed by different dynamics factors including interaction types (e.g., elastic force or electrostatic force) and interaction strengths (e.g., different amounts of charges on charged objects), as illustrated in Figure 7. In this work, we term these learning scenarios as continual dynamics learning (CDL), and for the first time formally formulate the setting of CDL. In CDL, continually learning from systems with varying dynamics may overwrite a model’s knowledge encoded in the model weights and trigger the catastrophic forgetting problem. In other words, the model may only accurately predict the most recently observed system dynamics while failing on earlier systems. This phenomenon is empirically verified and reported in Section 4.6. Moreover, many real-world systems exhibit repeated dynamics, and mitigating the forgetting is actually crucial in a broad range of scenarios. For example, dynamics of many physics systems are controlled by environmental factors, e.g. temperatures (Huang et al., 2023). The values of these factors like temperatures typically oscillate within a certain range, therefore the dynamics of the systems will also repeat. This is also true in biological systems. Moreover, biological cellular systems also go through different phases of cell cycle. Catastrophic forgetting phenomenon has also been observed in other fields like computer vision (Van de Ven & Tolias, 2019; Wang et al., 2024) and graph learning (Zhang et al., 2024b), and different approaches have been proposed to facilitate the models in the continual learning setting. However, as demonstrated in our experiments (Section 4.3), most existing continual techniques, which are based on regularization or memory-replay (Yoon et al., 2017; Kirkpatrick et al., 2017; Lopez-Paz & Ranzato, 2017), are designed to accommodate the patterns of different tasks  [^1] In our work, a task refers to a system with a specific dynamics patterns. While in other fields, e.g. Computer Vision, a task could refer to certain categories of images. within one set of model weights and fail to effectively alleviate the forgetting issue in the context of dynamics learning, especially when the consecutive systems contain different number of objects and exhibit significantly different dynamics patterns.
@@
Targeting the challenge, we turn to the parameter-isolation based continual learning techniques, and propose a novel mode-switching graph ODE (MS-GODE) model, which can continually learn over varying system dynamics and automatically switch to the optimal mode during the test stage. MS-GODE consists of three major components: a prediction network serving as the backbone, a sub-network learning module for encoding the observed dynamics into masks, and a mode-switching module for switching the sub-network mode based on the observation. To support irregular and incomplete observational data in the practical scenario studied in our work, our backbone network is built based on LG-ODE model (Huang et al., 2020), which has a Variational AutoEncoder (VAE) (Kingma & Welling, 2013) structure facilitated by ODE-based prediction (Rubanova et al., 2019). The basic workflow of MS-GODE is as follow: Given the observational data, an encoder network first encodes the data into latent states, upon which an ODE-based generator predicts the future system trajectories within the latent space. Finally, a decoder network maps the predicted latent states back into the data space. Upon this framework , we adopt the sub-network learning strategy (Wortsman et al., 2020; Ramanujan et al., 2020; Zhou et al., 2019), which fixes the model weights after initialization and optimize a unique binary mask over the parameters for each system during training. In this way, unlike standard training strategy that encodes data patterns solely in one set of model weights, MS-GODE encodes different types of dynamics in different sub-networks by the collaboration between the binary masks and the fixed-weight backbone network. In the test stage, the model will be switched to the optimal mode by the switching module via selecting the the most suitable mask that can most accurately reconstruct the given observation. Moreover, targeting that the existing entropy-based mask selection technique is only applicable to classification tasks, we further develop the novel observation reconstruction based mask selection strategy in the mode-switching module. In this way, despite the significant difference between consecutive systems, catastrophic forgetting is avoided.
@@
Besides innovatively formulating the CDL setting and the technical contribution of an effective model in CDL scenario, we have also created a novel dynamic system benchmark, Bio-CDL consisting of biological cellular systems based on the VCell platform (Schaff et al., 1997; Cowan et al., 2012; Blinov et al., 2017). Compared to the widely adopted simulated physics systems, cellular systems contain heterogeneous objects and interactions, offering richer and more challenging patterns of system dynamics. Therefore, Bio-CDL will significantly enhance the research of machine learning-based system dynamics prediction. In experiments, we thoroughly investigate the influence of the system sequence configuration on model performance in CDL with both the widely adopted physics systems and our newly constructed BioCDL, which demonstrates the advantage of MS-GODE over existing state-of-the-art techniques.
Figure 2: A molecule system may enter different phases and exhibit different dynamics as environmental factors (e.g. temperature) change. Molecules in solid state can only vibrate at fixed locations because of the strong interaction between them. Upon entering the liquid state, the interaction strength decreases and molecules can move around. In gas state, molecules move more freely with little molecule-wise interaction.
@@ 2.1 Learning based dynamic System Prediction
In recent years, graph neural networks (GNNs) have been proven to be promising in modeling and predicting the complex evolution of systems consisting of interacting objects (Battaglia et al., 2016; Kipf et al., 2018; Sanchez-Gonzalez et al., 2019; Huang et al., 2022; Liu et al., 2024; Luo et al., 2024b; Zhang et al., 2019; Luo et al., 2024a; Wu et al., 2021; Luo et al., 2025; Jiang et al., 2024; Guo et al., 2024; Zhang et al., 2020b; Shen et al., 2024; Yin et al., 2024; Chen et al., 2025). This was firstly demonstrated by (Battaglia et al., 2016) with Interaction Network (IN), which iteratively infers the effects of the pair-wise interactions within a system and predicts the changes of the system states. Following IN, (Kipf et al., 2018) proposed neural relational inference (NRI) to predict systems consisting of objects with unknown relationships. (Mrowca et al., 2018) proposed the Hierarchical Relation Network (HRN) that extends the predictions to systems consisting of deformable objects. (Sanchez-Gonzalez et al., 2019) proposed the Hamiltonian ODE graph network (HOGN), which injects Hamiltonian mechanics into the model as a physically informed inductive bias. Later, to better consider the intrinsic symmetry of the target systems, GNNs with different invariance and equivariance are proposed (Satorras et al., 2021; Huang et al., 2022; Han et al., 2022; Brandstetter et al., 2021; Wu et al., 2023; Liu et al., 2024). To better capture the complex system interactions, High-order graph ODE (HOPE) (Luo et al., 2023) innovatively incorporates information from high-order spatial neighborhood and high-order derivatives into dynamical system modeling. Considering that the observation of real-world systems may be incomplete and irregular samples, (Huang et al., 2020) proposed LG-ODE, which is capable of generating continuous system dynamics based on the latent ordinary differential equations. Later, Coupled Graph ODE (CG-ODE) (Huang et al., 2021) was proposed to apply ODE-based modeling to both node features and interactions. Similar ideas are also adopted in other time series research (Rubanova et al., 2019). By separating the commonalities intrinsic to the systems and the environmental factors causing the dynamics shift, Generalized Graph Ordinary Differential Equations (GG-ODE) (Huang et al., 2023) improves the generalization across systems in different environments. Similarly, Prototypical Graph ODE (PGODE) (Luo et al., ) disentangles object states and system states to independently model their influence and improve the generalization capability. Disentangled Intervention-based Dynamic graph Attention networks (DIDA) (Zhang et al., 2022c) disentangles the invariant and variant patterns in dynamic graphs, and leverages the invariant patterns to ensure a stable prediction performance under spatio-temporal distribution shift. Context-attended Graph ODE (CARE) (Luo et al., 2024c) models the continuously varying environmental factors with a context variable, which is leveraged to better predict the system evolution with temporal environmental variation. Online Relational Inference (ORI) (Kang et al., 2024) models the relationship as trainable parameters, which is accompanied by AdaRelation for the online setting. Despite the substantial contributions these methods have made to dynamic system prediction, they have been limited to learning a single system with fixed dynamics. As one of the two major components of our MS-GODE, the backbone model for dynamics system prediction is mainly built upon the LG-ODE framework (Huang et al., 2020). Different from the other approaches that are typically limited to data with regular intervals or complete observation at each time stamp, LG-ODE supports irregular and incomplete observations, which is essential to the applications studied in our work.
@@ 2.2 Continual Learning & Masked Networks
Existing continual learning methods can be categorized into three types (Parisi et al., 2019; Van de Ven & Tolias, 2019; De Lange et al., 2021; Zhang et al., 2022a; Konishi et al., 2023; Zhang et al., 2024a; 2023b; 2022b; Fionda et al., 2023). Regularization-based methods slow down the adaption of important model parameters via regularization terms, so that the forgetting problem is alleviated (Wu et al., 2024; Goswami et al., 2023). For example, Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017) and Memory Aware Synapses (MAS) (Aljundi et al., 2018) estimate the importance of the model parameters to the learned tasks, and add penalty terms to slow down the update rate of the parameters that are important to the previously learned tasks. Second, experience replay-based methods replay the representative data stored from previous tasks to the model when learning new tasks to prevent forgetting (Liang & Li, 2023; Rolnick et al., 2019; Rebuffi et al., 2017; Prabhu et al., 2020). For example, Gradient Episodic Memory (GEM) (Lopez-Paz & Ranzato, 2017) leverages the gradients computed based on the buffered data to modify the gradients for learning the current task and avoid the negative interference between learning different tasks. Finally, parameter isolation-based methods gradually introduce new parameters to the model for new tasks to prevent the parameters that are important to previous tasks (Qiao et al., 2023; Yoon et al., 2017). For example, Progressive Neural Network (PNN) (Rusu et al., 2016) allocates new network branches for new tasks, such that the learning on new tasks does not modify the parameters encoding knowledge of the old tasks. Our proposed MS-GODE also belongs to the parameter isolation-based methods, and is related to subnetwork-based ones (Wortsman et al., 2020; Kang et al., 2022; Zhou et al., 2019) and the edge-popup algorithm (Ramanujan et al., 2020). SupSup studies continual learning for classification tasks with an output entropy-based mask selection, which is not applicable to our task. Edge-popup algorithm provides a simple yet efficient strategy to select a sub-network, and is adopted by us to optimize the binary masks over the model parameters.
@@ 3.1 Preliminaries
In CDL, a model is required to sequentially learn on multiple systems. A system is composed of multiple interacting objects, and is naturally represented as a graph $\mathcal{G}=\{\mathbb{V},\mathbb{E}\}$. $\{\mathbb{V}$ is the node set denoting the objects of the system, and $\mathbb{E}$ is the edge set containing the information of the relation and interaction between the objects. Based on $\mathbb{E}$, the spatial neighbors of a node $v$ is defined as $\mathcal{N}_{s}(v)=\{u|e_{u,v}\in\mathbb{E}\}$. Each object node $v$ is accompanied by observational data containing the observed states at certain time steps $\mathbb{X}_{v}=\{\mathbf{x}_{v}^{t}|t\in\mathbb{T}_{v}\}$, i.e. the trajectory of the system evolution. With a system structured as a graph, its trajectory is naturally a spatial-temporal graph, in which each node is an observed state $\mathbf{x}_{v}^{t}$. In the following, we will refer to $\mathbf{x}_{v}^{t}$ as a ‘state’. In a multi-body system, the trajectory records the 3D locations of the particles over time. While in other systems, e.g. cellular systems, a state could be the amount of a certain substance instead of locations, and a trajectory records the states at different time stamps. The set $\mathbb{T}_{v}$ contains the time steps (real numbers) when the states of $v$ are observed and can vary across different objects. For the prediction task, the observations lie within a certain period, i.e. $\bigcup_{v\in\mathbb{V}}\mathbb{T}_{v}\in$[$t_{0},t_{1}$], and the task is to predict the system states at future time steps beyond $t_{1}$. We denote the future time steps to predict for an object $v$ as $\mathbb{T}_{v}^{pred}$. In CDL, different systems in a sequence may exhibit different dynamics and contain different objects (i.e. $\mathbb{V}$).
Figure 3: The upper part illustrates the workflow of MS-GODE during mask selection and inference, which are denoted by dashed and solid lines. During mask selection, the observation is split into two parts, and the first part is fed into the model for selecting the mask that can best predict the second part. During inference, the entire observation is fed into the model for prediction of future states. The lower part demonstrates the structure of the masked encoder, masked generator, and masked decoder. Different components fetch their corresponding mask from the mask pool and apply the mask onto the parameters.
@@ 3.2 Framework Overview
In this subsection, we provide a high-level introduction to the workflow of MS-GODE (Figure 3), while the details of each component are provided in the following subsections.
Overall, MS-GODE consists of three core components: 1) The backbone network; 2) The sub-network learning module; 3) The mode switching module. The backbone network consists of: 1) an encoder network $\mathrm{Enc}(\cdot;\theta_{E})$ parameterized by the parameters $\theta_{E}$ for encoding the trajectories into the latent space; 2) An ODE-based generator $\mathrm{Gen}(\cdot;\theta_{G})$ parameterized by $\theta_{G}$ for predicting the future trajectories within the latent space; 3) A decoder network $\mathrm{Dec}(\cdot;\theta_{D})$ parameterized by $\theta_{D}$ for transforming the predicted latent states back into the data space. Within a standard learning scheme, the model will be trained by updating the parameters $\{\theta_{E},\theta_{G},\theta_{D}\}$. However, as introduced above, when continually training the model on multiple systems with different dynamics, this learning scheme will bias the model towards the most recently observed system, causing the catastrophic forgetting problem. In this work, inspired by the recent advances in sub-network learning (Ramanujan et al., 2020; Wortsman et al., 2020), we propose to avoid direct optimization of the parameters $\{\theta_{E},\theta_{G},\theta_{D}\}$. Instead, we train the model by optimizing the connection topology of the backbone model and encode the dynamics of each system into a sub-network. This is equivalent to encoding each system-specific dynamics pattern into a binary mask overlaying the shared parameters with fixed values. In this way, the interference between learning on different systems with different dynamics, i.e. the forgetting problem, can be avoided. Moreover, this approach is also memory efficient since the space for storing the binary masks is negligible. With the system-specific masks, the three model components are formulated as: